OpenStudy (anonymous):

An orchestra conductor selects the first, second, and third chair violinists for an upcoming performance. If the conductor can choose from 210 combinations of first, second, and third chair violinists, how many violinists are in the orchestra.

OpenStudy (anonymous):

A little confused with permutation and combinations.

OpenStudy (anonymous):

SAT question and not for marks just fyi.

OpenStudy (anonymous):

this question is kind of asking it backwards what is \(n\) so that \[_nP_3=120\]

OpenStudy (anonymous):

How do I solve it from there tho?

OpenStudy (anonymous):

try some numbers

OpenStudy (anonymous):

I know the answer is 7 just don't know the steps.

OpenStudy (anonymous):

\[10\times 9\times 8=720\] way too big

OpenStudy (anonymous):

guess and check \[6\times 5\times 4=120\] too small

OpenStudy (anonymous):

oh no it is not too small it is just right! so i would say the answer is 6, not 7

OpenStudy (anonymous):

oops 210 not 120 ok go bigger